The Perimeter of a rectangle is 52 feet. The length and width are represented by two consecutive even integers. Find the dimensions of the rectangle.
Answer by checkley77(12844)

Let x & x+2 be the Length & Width.
2L+2W=52
2X+2(X+2)=52
2X+2X+4=52
4X=52-4
4X=48
X=48/4
X=12 ANS. FOR THE LENGTH.
12+2=14 ANS. FOR THE WIDTH.
PROOF:
2*12+2*14=52
24+28=52
52=52